# Gram-like matrix preserving extensions of noncommutative polynomials to sum of Hermitian Squares

### v1

@@ 3.2. Block Matrix Technique
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative SOHS polynomial. Then $f+c$ is again a SOHS polynomial for any $c>0$.
@@
    If $f$ has a constant term, there there exists a Gram-like matrix of $f+c$ that can be obtained from a Gram-like matrix of $f$ by changing only one diagonal entry.
@@
    If $f$ does not have a constant term, then there exists a Gram-like matrix of $f+c$ that contains a Gram-like matrix of $f$ as a leading principal submatrix.
@@
    As $f$ is a SOHS of degree $2d$ (say), and it has a constant term, then by Theorem 3.5 there exists a positive semidefinite Gram-like matrix $G_{f}$ and a column vector $W_{k}$ of monomials of degree at most $d$ with $1$ as the $(i_{0},1)$-th entry such that $f=W_{k}^{\ast}\;G_{f}\;W_{k}$. We now consider the matrix $\widetilde{G_{f}}$, given as follows :
    $(\widetilde{G_{f}})_{ij}=\left\{\begin{array}[]{ll}(G_{f}){{}_{ij}}&\mbox{if $(i,j)\neq(i_{0},i_{0})$};\\ \\ (G_{f}){{}_{i_{0}i_{0}}}+c&\mbox{\text{otherwise}}.\end{array}\right.$ (12)
    Then clearly, as $c>0$, all the principal minors of $\widetilde{G_{f}}$ is greater than equal to that of $G_{f}$ and $G_{f}$ is a positive semidefinite matrix. Therefore, $\widetilde{G_{f}}$ is also so by Theorem 2.5. Moreover,
    $f+c=W_{k}^{\ast}\;\widetilde{G_{f}}\;W_{k},$ (13)
    where $\widetilde{G_{f}}$ is as in (12). So, the assertion follows from (13) and Theorem 3.5.
@@
    As $f$ is a SOHS of degree $2d$ (say), and it is not having any constant term, by Theorem 3.5 there exists a positive semidefinite Gram-like matrix $G_{f}$ and a column vector $W_{k}$ of monomials of degree at most $d$ with none of its entries being $1$ such that $f=W_{k}^{\ast}\;G_{f}\;W_{k}$. We now consider the matrix $\widetilde{G_{f}}$ of size $k+1$, given as follows :
    $\widetilde{G_{f}}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{f}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}c\end{matrix}\end{array}\right)\;.$ (14)
    We also define $\widetilde{W_{k+2}}$ as follows.
    $\widetilde{W_{k+1}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}\end{matrix}\\ \begin{matrix}1\end{matrix}\end{array}\right)\;.$ (15)
    Then, clearly we have
    $f+c\;=\;\widetilde{W_{k+1}}^{\ast}\;\widetilde{G_{f}}\;\widetilde{W_{k+1}},$ (16)
    where $\widetilde{W_{k+1}}$ and $\widetilde{G_{f}}$ are as in (15) and (14) respectively. Now as $c>0$ and $G_{f}$ is a positive semidefinite matrix, so is $\widetilde{G_{f}}$ by Lemma 2.7. Hence, the assertion follows from (16) and Theorem 3.5.
Corollary 3.11.
Let $A$ be a positive semidefinite matrix of size $n$. Then for any positive real numbers $b_{1},b_{2},\cdots b_{n}$, the matrix $\widetilde{A}$ given by
$\widetilde{A}_{ij}=\left\{\begin{array}[]{ll}A_{ij}&\mbox{if $i\neq j$};\\ \\ A_{ii}+b_{i}&\mbox{\text{otherwise}}.\end{array}\right.$
Proof.
Follows from the repeated applications of proof of the part $(1)$ of Theorem 3.10 and Lemma 2.6. ∎
Example 3.12.
Consider $f=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}\in\mathbb{R}\langle X_{1},X_{2}\rangle$. Clearly, $f$ is not a SOHS as it doesn’t even lie in ${\rm Sym}\;\mathbb{R}\langle X_{1},X_{2}\rangle$. Now we consider the polynomials $\widetilde{f}_{i}$ s, by adding some terms to $f$, given as follows :
$\begin{split}\widetilde{f}_{1}&=f+3X_{2}X_{1}+4X_{2}X_{1}^{2}X_{2}+1=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+4X_{2}X_{1}^{2}X_{2}+1,\\ \widetilde{f}_{2}&=f+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+4=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+4,\\ \widetilde{f}_{3}&=f+3X_{2}X_{1}+9X_{2}X_{1}^{2}X_{2}+1=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+9X_{2}X_{1}^{2}X_{2}+1,\\ \widetilde{f}_{4}&=f+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+9=X_{1}^{2}+X_{2}^{2}+3X_{1}X_{2}+3X_{2}X_{1}+X_{2}X_{1}^{2}X_{2}+9.\end{split}$
The SOHS extensions $\widetilde{f}_{1}$ and $\widetilde{f}_{2}$ of $f$ can be decomposed as follows:
$\begin{split}\widetilde{f}_{1}&=(X_{1}+X_{2})^{\ast}(X_{1}+X_{2})+(2X_{1}X_{2}+1)^{\ast}(2X_{1}X_{2}+1),\\ \widetilde{f}_{2}&=(X_{1}+X_{2})^{\ast}(X_{1}+X_{2})+(X_{1}X_{2}+2)^{\ast}(X_{1}X_{2}+2).\end{split}$ (17)
Similarly, the SOHS extensions $\widetilde{f}_{3}$ and $\widetilde{f}_{4}$ of $f$ has following decompositions:
$\begin{split}\widetilde{f}_{3}&=X_{1}^{\ast}X_{1}+X_{2}^{\ast}X_{2}+(3X_{1}X_{2}+1)^{\ast}(3X_{1}X_{2}+1),\\ \widetilde{f}_{4}&=X_{1}^{\ast}X_{1}+X_{2}^{\ast}X_{2}+(X_{1}X_{2}+3)^{\ast}(X_{1}X_{2}+3).\end{split}$ (18)
@@
    It can be noted that in the first set of decompositions, as given in (17), pre existing SOHS part of $f$ is getting affected. In contrast, in the decompositions, as given in (18), the same is remaining intact.
@@
    The decompositions, as in (18), are more natural in the sense that these do not depend on any sort of relation between the SOHS part and non-SOHS part of $f$, and therefore have the potential to provide positive semidefinite Gram-like matrices of $\widetilde{f}_{i}$ s, $i=3,4$, incorporating positive semidefinite Gram-like matrices of $f$.
We record these observations in a proper setup through the following theorems.
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial which can be expressed as a SOHS plus $n$ many terms. Then it is always possible to obtain a SOHS $\widetilde{f}$ by adding at most $12n+$ 1 many terms with $f$.
@@
Let ${\rm deg}(f)=d$. By hypothesis, $f$ can be written as follows:
$f=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{n}a_{j}\zeta_{j},$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $1\leq i\leq m$, and $1\leq{\rm deg}(\zeta_{j})\leq d$ for all $1\leq j\leq n$. We define a polynomial $h$, consisting of $12n+$ 1 terms, as follows:
$h:=\sum_{j=1}^{n}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{n}\zeta_{j}^{\ast}\zeta_{j}+(\sum_{j=1}^{n}a_{j}^{2}).$
Then $\widetilde{f}(:=f+h)$ is a SOHS as it has the following decomposition:
$\widetilde{f}=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{n}(\zeta_{j}+a_{j})^{\ast}(\zeta_{j}+a_{j}).$
We now aim to provide a positive semidefinite Gram-like matrix for the extended SOHS $\widetilde{f}$ in terms of a positive semidefinite Gram-like matrix of the partial SOHS $f$. For that we put some conditions on the non-SOHS part of $f$ and then use block-matrix technique, as introduced in the proof of Part - 2 of Theorem 3.10. We first prove it assuming that nonconstant part of $f$ is made out of a single term.
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, which can be expressed as a SOHS plus another term $a\zeta$. Assume that ${\rm rc}(\eta,i)\neq\zeta$ for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$. Then
@@
    It is always possible to obtain a SOHS $\widetilde{f}$ by adding at most two nonconstant terms and a constant term with $f$ such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of the SOHS part of $f$ as a leading principal submatrix.
  2. (2)
    In fact, infinitely many such $\widetilde{f}$ exist.
Proof.
Let ${\rm deg}(f)=d$. By hypothesis, $f$ can be written as follows:
$f=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+a\zeta,$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $i$ and $1\leq{\rm deg}(\zeta)\leq d$. Let $G$, say of order $k$, be a positive semidefinite Gram-like matrix for the SOHS part $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ of $f$ such that there exist a column vector $W_{k}$ satisfying
  1. (a)
    $1$ does not appear as an entry of $W_{k}$,
  2. (b)
    $\zeta$ does not appear as an entry of $W_{k}$. and
  3. (c)
    $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}=W_{k}^{\ast}\;G\;W_{k}.$ (19)
It can be noted that existence of such $W_{k}$ satisfying
  • (b) is guaranteed by the hypothesis ${\rm rc}(\eta,i)\neq\zeta$ for any $i\geq 0$ for any monomial $\eta$ of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ (Indeed, let $\zeta$ appears as an entry of $W_{k}$. Then $\zeta^{\ast}\zeta$ appears in the expansion of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ as its coefficient in nonzero by (9b) of Theorem 3.5. But then ${\rm rc}(\zeta^{*}\zeta,{\rm deg}(\zeta))=\zeta$.),
  • (c) is guaranteed by Theorem 3.5.
Let us choose any $b,c\in\mathbb{R}$ such that
$\begin{split}b+c&>0,\text{\;and\;}\\ bc-a^{2}&>0.\end{split}$ (20)
Now we consider the matrix $\widetilde{G}$ given as follows:
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}b&a\\ a&c\end{matrix}\end{array}\right)\;.$ (21)
We also define $\widetilde{W_{k+2}}$ as follows.
$\widetilde{W_{k+2}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}^{t}&1&\zeta\end{matrix}\end{array}\right)^{t}\;.$ (22)
Then we have
$\widetilde{f}=f+a\;\zeta^{\ast}+c\;\zeta^{\ast}\zeta+b=\widetilde{W_{k+2}}^{\ast}\;\widetilde{G}\;\widetilde{W_{k+2}},$ (23)
where $\widetilde{G}$ and $\widetilde{W_{k+2}}$ are as in (21) and (22) respectively. Now as the matrices $G$, as in (19), and $\begin{pmatrix}b&a\\ a&c\\ \end{pmatrix}$, due to (20), are positive semidefinite matrices, so is the matrix $\widetilde{G}$ by Lemma 2.7. Hence $\widetilde{f}$, as in (23), is a SOHS by Theorem 3.5 and clearly obtained from $f$ by adding two nonconstant monomials $a\;\zeta^{\ast}$ and $c\;\zeta^{\ast}\zeta$ and a constant term $b$. Moreover if $\zeta^{\ast}=\zeta$, then at least the term $a\zeta^{\ast}$ is not needed in $\widetilde{f}$ In fact, in that case we can take $\widetilde{G}$ as
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}b&\tfrac{a}{2}\\ \tfrac{a}{2}&c\end{matrix}\end{array}\right)\;,$
where $b+c>0$ and $04bc-a^{2}>$ 0.
As there are infinitely many choices for $b$ and $c$ satisfying (20), we have infinitely many choices for $\widetilde{G}$ and hence for $\widetilde{f}$. ∎
We now extend Theorem 3.15 when non-SOHS part of $f$ is having multiple terms.
Theorem 3.16.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, which can be expressed as a SOHS plus $r$ many terms $a_{1}\zeta_{1},a_{2}\zeta_{2},\ldots,a_{r}\zeta_{r}$. Assume that ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$. Then
  1. (1)
    It is always possible to obtain a SOHS $\widetilde{f}$ by adding at most $12r+$ 1 terms with $f$ such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of the SOHS part of $f$ as a leading principal submatrix.
@@
$f=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}a_{j}\zeta_{j},$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $1\leq i\leq m$ and $1\leq{\rm deg}(\zeta_{j})\leq d$ for all $1\leq j\leq r$. Let $G$, say of order $k$, be a positive semidefinite Gram-like matrix for the SOHS part $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ of $f$ such that there exist a column vector $W_{k}$ satisfying
@@
  • (b) is guaranteed by the hypothesis ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$ (Indeed, let $\zeta_{j}$ appears as an entry of $W_{k}$ for some $j$. Then $\zeta_{j}^{\ast}\zeta_{j}$ appears in the expansion of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ as its coefficient in nonzero by (9b) of Theorem 3.5. But then ${\rm rc}(\zeta_{j}^{\ast}\zeta_{j},{\rm deg}(\zeta_{j}))=\zeta_{j}$.),

### v2

@@ 3.2. Gram-like matrix preserving SOHS extensions
Let $f=h+\sum_{j=1}^{m}d_{j}\delta_{j}\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial with a given SOHS part $h$ and the monomials $\delta_{j}$s are not Hermitian squares.
@@
    Assume that there exists a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ w.r.t $h$. Then, there exists a representation
    $\mathbf{R}_{h}\;:\;h=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;,$
    and, for all $1\leq j\leq m$, there exist rc decompositions of $\delta_{j}=\beta_{n_{j}}^{\ast}\gamma_{n_{j}}$, for some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that no $\beta_{n_{j}}$ and $\gamma_{n_{j}}$ simultaneously lie in $\{\zeta_{1},\ldots,\zeta_{k}\}$.
@@
    Conversely, assume that there exists a representation
    $\mathbf{R}_{h}\;:\;h=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;,$
    and, for all $1\leq j\leq m$, there exist rc decompositions of $\delta_{j}=\beta_{n_{j}}^{\ast}\gamma_{n_{j}}$, for some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that no $\beta_{n_{j}}$ and $\gamma_{n_{j}}$ simultaneously lie in $\{\zeta_{1},\ldots,\zeta_{k}\}$. Then, there exists a partially specified symmetric matrix $G$ of the form
    $G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;,$
    where $G_{h}$ is the Gram-like matrix of $h$ w.r.t the representation $\mathbf{R}_{h}$. Moreover, if $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$, then there exist infinitely many Gram-like matrix preserving extensions of $f$.
@@
    As $\widetilde{f}$ is a Gram-like matrix preserving extension w.r.t $h$, by Lemma 3.14 it must be of the form
    $G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;$
    up to a permutation, where $G_{h}$ is a positive semidefinite Gram-like matrix of $h$. Let $G_{h}$ be the Gram-like matrix associated with
    $\mathbf{R}_{h}\;:\;h=\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;.$
    1. Case 1
      All the coefficients $d_{j}$ of the monomials $\delta_{j}$ of $f-h$ are distributed in $B$ :
      In this case, there exist some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that the decompositions
      $\delta_{j}=({\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j}))^{\ast}{\rm rc}(\delta_{j},n_{j})$
      satisfies
      ${\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j})\neq\zeta_{i}\neq{\rm rc}(\delta_{j},n_{j}),$
      for all $1\leq i\leq k$ and $1\leq j\leq m$.
    2. Case 2
      Not all the coefficients $d_{j}$ of the monomials $\delta_{j}$ of $f-h$ are distributed in $B$ :
      In this case, some $d_{j}$ must be distributed in $A$ or $A^{t}$ or both. Therefore,
      1. (i)
        either there exist decompositions $\delta_{j}=({\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j}))^{\ast}{\rm rc}(\delta_{j},n_{j})$, for some $0\leq n_{j}\leq{\rm deg}(\delta_{j})$, such that
        $\beta_{n_{j}}(:={\rm rc}(\delta_{j}^{\ast},{\rm deg}(\delta_{j})-n_{j}))=\zeta_{i}$
        for some $i\in K_{1}\subseteq\{1,\ldots,k\}$ and $j\in M_{1}\subseteq\{1,\ldots,m\}$. Then, we must have
[…]
        as otherwise these $d_{j}$s would have been distributed inside $G_{h}$, which is not possible as $\widetilde{f}$ is a Gram-like matrix preserving extension of $f$ w.r.t $h$.
      2. (ii)
        Exactly similar to (i).
@@ 2. Case 2
    Reversing the implications of the proof the first part of the assertion, it is easy to see that if no $\beta_{n_{j}}$ and $\gamma_{n_{j}}$ simultaneously lie in $\{\zeta_{1},\ldots,\zeta_{k}\}$, then there exists a partially specified symmetric matrix $G$ of the form
    $G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;,$
    where $G_{h}$ is the Gram-like matrix of $h$ w.r.t the representation $\mathbf{R}_{h}$, such that the coefficient $d_{j}$s of $\delta_{j}$s and $\delta_{j}^{\ast}$s are distributed in $G$ outside the block $G_{h}$. Here we are taking $W_{l}$ as a associated monomial vector to $G$ obtained by taking $\zeta_{i}$s as the first $k$ constituent monomials and taking $\beta_{n_{j}}$s and $\gamma_{n_{j}}$s (without any repetition) as the remaining monomials. Also $d_{j}$s can’t be distributed in the diagonals of $B$, as $\delta_{j}$s are not Hermitian squares.
    Therefore we now have a partially specified symmetric matrix $G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)$ such that $G_{h}$ and $B$ are square matrices and the only unspecified entries of $G$ are the diagonal entries of $B$. Now, if $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$, then by Lemma 3.17, we can choose diagonal entries of $B$ in such a way that makes $G$ a positive semidefinite matrix. Therefore, the SOHS $\widetilde{f}:=W_{l}^{\ast}\;G\;W_{l}$ is the required Gram-like preserving extension of $f$.
    Moreover, it is clear from Lemma 3.17 that there are infinitely many choices for the diagonal entries of $B$ that makes $G$ positive semidefinite. Therefore, we obtain infinitely many such Gram-like preserving extension $\widetilde{f}$ of $f$.
@@
    In Theorem 3.19, the hypothesis that the monomials $\delta_{j}$s of $f-h$ are non Hermitian is used to prove the converse part only.
@@
    Whenever the matrix $G_{h}$ of $h$, as in Theorem 3.19, is positive definite, then the necessary condition itself becomes sufficient as the condition $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$ is satisfied by default.
  3. (3)
[…]
    Assume that there exists a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ w.r.t $h$. Because of the possible rc decompositions of $X_{1}$ and $X_{2}$, as mentioned earlier, $\widetilde{f}$ must have a constant term, say $d$. Then for any such $\widetilde{f}$, there should be some positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ such that it contains either $\begin{pmatrix}1&1&1\\ 1&1&-1\\ 1&-1&d\end{pmatrix}$ or $\begin{pmatrix}1&1&-1\\ 1&1&1\\ -1&1&d\end{pmatrix}$, corresponding to monomial vector $\begin{pmatrix}X_{1}&X_{2}&1\end{pmatrix}^{t}$ or $\begin{pmatrix}X_{2}&X_{1}&1\end{pmatrix}^{t}$, as a principal minor. But any of these principal minors are having determinant $-4(<0)$. Therefore, by Lemma 2.7, such $G_{\widetilde{f}}$, and hence such $\widetilde{f}$, can’t exist.
Example 3.21.
[…]
Now we want to look that under what circumstances, a Gram-like matrix preserving extension $\widetilde{f}$ of $f=h+\sum_{j=1}^{m}d_{j}\delta_{j}$ admits a Gram-like matrix $G_{\widetilde{f}}$ of the form $G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)$. Here we omit the additional assumption that the monomials $\delta_{j}$s are not Hermitian squares.
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial. Let $h$ be a SOHS part of $f$. Let $\widetilde{f}$ be a SOHS extension of $f$. Then the following are equivalent :
  1. (1)
    $\widetilde{f}$ is a Gram-like matrix preserving extension of $f$ w.r.t $h$ such that there exists a positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ of $\widetilde{f}$ of the form
    $G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)\;,$ (11)
    where $G_{h}$ and $G_{\widetilde{f}-h}$ are $k\times k$ and $l\times l$ positive semidefinite Gram-like matrices of $h$ and $\widetilde{f}-h$ respectively.
  2. (2)
    There exist a representation $\mathbf{R}_{h}$ of $h$ and and a representation $\mathbf{R}_{\widetilde{f}-h}$ of $\widetilde{f}-h$, given as
    $\begin{split}\mathbf{R}_{h}\;:\;h=&\sum_{i=1}^{k}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq k}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i})\;,\\ \mathbf{R}_{\widetilde{f}-h}\;:\;\widetilde{f}-h=&\sum_{i=1}^{l}c_{i}\eta_{i}^{\ast}\eta_{i}+\sum_{1\leq i<j\leq l}d_{ij}(\eta_{i}^{\ast}\eta_{j}+\eta_{j}^{\ast}\eta_{i}),\end{split}$
    both satisfying conditions as in (7), such that
    $\zeta_{i}\neq\eta_{j}\;\text{for\;all\;}1\leq i\leq k,1\leq j\leq l.$
@@
$(1)\Rightarrow(2)$ : By Definition 3.11 and Remark 3.12, there exists a monomial vector
$\widetilde{W_{k+l}}:=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}&\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}$
associated to $G_{\widetilde{f}}$ such that $h=W_{k}^{\ast}\;G_{h}\;W_{k}$. Here $W_{k}=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}\;^{t}$. Then,
$\begin{split}\widetilde{f}-h&=\widetilde{W_{k+l}}^{\ast}\;\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)\;\widetilde{W_{k+l}}-W_{k}^{\ast}\;G_{h}\;W_{k}\\ &=W_{l}^{\ast}\;G_{\widetilde{f}-h}\;W_{l}\;,\end{split}$
where $W_{l}=\begin{pmatrix}\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}$. Now, let $\mathbf{R}_{h}$ and $\mathbf{R}_{\widetilde{f}-h}$ be the associated representations of $G_{h}$ and $G_{\widetilde{f}-h}$ respectively. Then, $\mathbf{R}_{h}$ and $\mathbf{R}_{\widetilde{f}-h}$ must satisfy the conditions as in (7). Moreover, $\zeta_{i}\neq\eta_{j}$ for all $1\leq i\leq k$, $1\leq j\leq l$, as $\widetilde{W_{k+l}}$ is a monomial vector.
$(2)\Rightarrow(1)$ : Let $G_{h}$ and $G_{\widetilde{f}-h}$ be positive semidefinite Gram-like matrices of $h$ and $\widetilde{f}-h$ associated with representations $\mathbf{R}_{h}$ and $\mathbf{R}_{\widetilde{f}-h}$ respectively. As, $\zeta_{i}\neq\eta_{j}$ for all $1\leq i\leq k$, $1\leq j\leq l$, we can form the the monomial vector
$\widetilde{W_{k+l}}:=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}&\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}.$
Then, denoting $\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}\end{pmatrix}\;^{t}$ and $\begin{pmatrix}\eta_{1}&\cdots&\eta_{l}\end{pmatrix}\;^{t}$ by $W_{k}$ and $W_{l}$ respectively, we have:
$\begin{split}\widetilde{f}&=h+(\widetilde{f}-h)\\ &=W_{k}^{\ast}\;G_{h}\;W_{k}+W_{l}^{\ast}\;G_{\widetilde{f}-h}\;W_{l}\\ &=\widetilde{W_{k+l}}^{\ast}\;\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}G_{\widetilde{f}-h}\end{matrix}\end{array}\right)\;\widetilde{W_{k+l}}\;.\end{split}$
Remark 3.23.
Although the sum $h_{1}+h_{2}$ of two SOHS polynomials $h_{1}$ and $h_{2}$ is always SOHS, the difference $h_{1}-h_{2}$ does not enjoy the same. An example would be $h_{1}=1+X_{1}^{2}+X_{1}X_{2}+X_{2}X_{1}+X_{2}^{2}$ and $h_{2}=1+X_{1}^{2}$, as $h_{1}-h_{2}=X_{1}X_{2}+X_{2}X_{1}+X_{2}^{2}$ is not a SOHS. Comparing this with Example 3.21, we conclude that an important application of Theorem 3.22 is to produce nontrivial SOHS polynomials such that their difference is also a SOHS.
Theorem 3.22 talks about an existential criterion for a Gram-like matrix $G_{\widetilde{f}}$ of the extension $\widetilde{f}$ of $f$ of the form as in (11). We now present one such method of construction of $G_{\widetilde{f}}$ for a class of noncommutative polynomials.
@@
Let $f=h+\sum_{j=1}^{r}a_{j}\zeta_{j}\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, where $h$ is a SOHS. Assume that ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of $h$. Then
@@
    It is always possible to obtain a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ w.r.t $h$ by adding at most $12r+$ 1 terms with $f$.
@@
$f=h+\sum_{j=1}^{r}a_{j}\zeta_{j}=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}a_{j}\zeta_{j},$
where $a\in\mathbb{R}$, $g_{i}$s and $\zeta$ lie in $\mathbb{R}\langle\underline{X}\rangle$, ${\rm deg}(g_{i})\leq\tfrac{d}{2}$ if $d$ is even or ${\rm deg}(g_{i})\leq\tfrac{d-1}{2}$ if $d$ is odd, for all $1\leq i\leq m$ and $1\leq{\rm deg}(\zeta_{j})\leq d$ for all $1\leq j\leq r$. Let $G_{h}$, say of order $k$, be a positive semidefinite Gram-like matrix of $h$ such that there exist a column vector $W_{k}$ satisfying
@@
  • (b) is guaranteed by the hypothesis ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$ (Indeed, let $\zeta_{j}$ appears as an entry of $W_{k}$ for some $j$. Then $\zeta_{j}^{\ast}\zeta_{j}$ appears in the expansion of $\sum_{i=1}^{m}g_{i}^{\ast}g_{i}$ as its coefficient in nonzero by (7c) of Theorem 3.5. But then ${\rm rc}(\zeta_{j}^{\ast}\zeta_{j},{\rm deg}(\zeta_{j}))=\zeta_{j}$.),